Determine the inverse of A, and show that A


−1A = I.


A =


(2 1 0


2 −1 1


3 −2 4)





1
Expert's answer
2022-06-13T17:50:25-0400
"A=\begin{pmatrix}
 2 & 1 & 0\\
 2 & -1 & 1\\ 
 3 & -2 & 4\\ 
\end{pmatrix}"

Augment the matrix with the identity matrix:


"\begin{pmatrix}
 2 & 1 & 0 & & 1 & 0 & 0\\
 2 & -1 & 1 & & 0 & 1 & 0\\ 
 3 & -2 & 4 & & 0 & 0 & 1\\ 
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
 1 & 1/2 & 0 & & 1/2 & 0 & 0\\
 2 & -1 & 1 & & 0 & 1 & 0\\ 
 3 & -2 & 4 & & 0 & 0 & 1\\ 
\end{pmatrix}"

"R_2=R_2-2R_1"


"\begin{pmatrix}
 1 & 1/2 & 0 & & 1/2 & 0 & 0\\
 0 & -2 & 1 & & -1 & 1 & 0\\ 
 3 & -2 & 4 & & 0 & 0 & 1\\ 
\end{pmatrix}"

"R_3=R_3-3R_1"


"\begin{pmatrix}
 1 & 1/2 & 0 & & 1/2 & 0 & 0\\
 0 & -2 & 1 & & -1 & 1 & 0\\ 
 0 & -7/2 & 4 & & -3/2 & 0 & 1\\ 
\end{pmatrix}"

"R_2=-R_2\/2"


"\begin{pmatrix}
 1 & 1/2 & 0 & & 1/2 & 0 & 0\\
 0 & 1 & -1/2 & & 1/2 & -1/2 & 0\\ 
 0 & -7/2 & 4 & & -3/2 & 0 & 1\\ 
\end{pmatrix}"

"R_1=R_1-R_2\/2"


"\begin{pmatrix}
 1 & 0 & 1/4 & & 1/4 & 1/4 & 0\\
 0 & 1 & -1/2 & & 1/2 & -1/2 & 0\\ 
 0 & -7/2 & 4 & & -3/2 & 0 & 1\\ 
\end{pmatrix}"

"R_3=R_3+7R_2\/2"


"\begin{pmatrix}
 1 & 0 & 1/4 & & 1/4 & 1/4 & 0\\
 0 & 1 & -1/2 & & 1/2 & -1/2 & 0\\ 
 0 & 0 & 9/4 & & 1/4 & -7/4 & 1\\ 
\end{pmatrix}"

"R_3=4R_3\/9"


"\begin{pmatrix}
 1 & 0 & 1/4 & & 1/4 & 1/4 & 0\\
 0 & 1 & -1/2 & & 1/2 & -1/2 & 0\\ 
 0 & 0 & 1 & & 1/9 & -7/9 & 4/9\\ 
\end{pmatrix}"

"R_1=R_1-R_3\/4"


"\begin{pmatrix}
 1 & 0 & 0 & & 2/9 & 4/9 & -1/9\\
 0 & 1 & -1/2 & & 1/2 & -1/2 & 0\\ 
 0 & 0 & 1 & & 1/9 & -7/9 & 4/9\\ 
\end{pmatrix}"

"R_2=R_2+R_3\/2"


"\begin{pmatrix}
 1 & 0 & 0 & & 2/9 & 4/9 & -1/9\\
 0 & 1 & 0 & & 5/9 & -8/9 & 2/9\\ 
 0 & 0 & 1 & & 1/9 & -7/9 & 4/9\\ 
\end{pmatrix}"

We are done. On the left is the identity matrix. On the right is the inverse matrix.


"A^{-1}=\begin{pmatrix}
 2/9 & 4/9 & -1/9\\
 5/9 & -8/9 & 2/9\\ 
 1/9 & -7/9 & 4/9\\ 
\end{pmatrix}"

Check


"A^{-1}A=\begin{pmatrix}
 2/9 & 4/9 & -1/9\\
 5/9 & -8/9 & 2/9\\ 
 1/9 & -7/9 & 4/9\\ 
\end{pmatrix}\begin{pmatrix}
 2 & 1 & 0\\
 2 & -1 & 1\\ 
 3 & -2 & 4\\ 
\end{pmatrix}"

"=\begin{pmatrix}
 \dfrac{4+8-3}{9} & \dfrac{2-4+2}{9} & \dfrac{0+4-4}{9}\\
\\
 \dfrac{10-16+6}{9} & \dfrac{5+8-4}{9} & \dfrac{0-8+8}{9}\\
\\
 \dfrac{2-14+12}{9} & \dfrac{1+7-8}{9} & \dfrac{0-7+16}{9}\\
\\
\end{pmatrix}"

"=\begin{pmatrix}
 1 & 0 & 0\\
 0 & 1 & 0\\ 
 0 & 0 & 1\\ 
\end{pmatrix}=I"